# Better Call CLAUSE: A Discrepancy Benchmark for Auditing LLMs Legal Reasoning Capabilities

### v1

@@ Abstract
Manan Roy Choudhury1∗, Adithya Chandramouli1∗, Mannan Anand1 thanks: The first three authors are recognized for their shared contribution to this work., Vivek Gupta1 1Arizona State University {mroycho1, achand71, mannan.anand, vgupt140}@asu.edu
@@ 1 Introduction
The integration of Large Language Models (LLMs) into high-stakes legal work has outpaced the development of benchmarks that can audit their reliability against nuanced contractual flaws. This gap is critical, as a single undetected defect can have catastrophic consequences. In Perini Corp. v. Greate Bay Hotel & Casino, Inc., for instance, the simple omission of a consequential damages waiver resulted in a \$14.5 million liability, a sum over twenty times the contract’s fee McElroy, Deutsch, Mulvaney & Carpenter, LLP (2021). State-of-the-art models, optimized for fluency, are not trained to detect such legally significant risks, particularly those arising from the absence of text. Existing evaluation methods fail to probe these fragile reasoning capabilities Hu et al. (2025); Mishra et al. (2025), creating an urgent need for a new class of benchmark to stress-test LLMs against the adversarial and context-aware challenges of legal practice. This research directly confronts the ongoing debate about AI’s readiness for legal integration. By evaluating modern LLMs against our CLAUSE benchmark, we aim to provide empirical data to help answer the fundamental question: are these models truly reliable enough for such high-stakes work?
@@ 2 Our CLAUSE Dataset
Our pipeline applies a unified perturbation framework to the contracts. We extract full text and structural metadata (clause boundaries, section headers, cross-references) to preserve document integrity during modification. Contextually appropriate contradiction locations are identified using prompts across ten perturbation types. Three complementary prompting strategies ensure legal plausibility: persona-based roles (compliance officers, legal editors), few-shot examples, and chain-of-thought reasoning.
Each perturbation produces a modified contract with structured JSON metadata containing original/altered text, perturbation type, document location, and model-generated explanations. Legal contradictions include law citations and statutory context; in-text contradictions include conflicting clause references and logical justifications. The unified process is applied to both datasets ensuring consistent output quality and metadata format, supporting downstream detection accuracy and legal reasoning analysis tasks.
@@
Table 1 presents comprehensive statistics for both components, showing the distribution of JSON files and perturbations across all ten categories.
The dataset exhibits several notable characteristics. First, each JSON file contains multiple perturbations, with CUAD averaging 2.73 perturbations per file and the NLI split averaging 3.96 perturbations per file. Second, the distribution across categories is relatively balanced, though some variation exists due to the inherent complexity of generating specific perturbation types for different contract structures. Third, the legal categories generally contain fewer perturbations than their inText counterparts, reflecting the additional constraints imposed by statutory validation requirements.
@@
To validate quality and legal accuracy, we first engaged a legal expert to confirm the real-world relevance of our perturbation categories. Subsequently, 3 NLP experts evaluated 25% of our dataset (13,275 perturbation instances) using structured binary classification to determine meaningful legal contradictions.
Our human verification demonstrates exceptional quality with 98.58% contradiction validation rate and strong inter-rater reliability shown in Table 2. The evaluation examined semantic significance, contradiction strength, and contextual coherence, ensuring our benchmark contains high-quality examples for testing LLM legal reasoning capabilities.
@@ 5 Experimental Evaluation
explanation_match. The LLM, acting as a legal analyst, receives the combined explanations and all relevant GT contexts. It returns a structured JSON object with scores between 0-5 for: Accuracy (factual correctness), Completeness (coverage of key points), Clarity (ease of understanding), Legal Reasoning (soundness of legal logic), and a final binary Match flag that determines the alignment outcome.
This process is adapted for two contradiction types. Ground-truth context are provided during the judging process for better context aware outputs: a) In-Text Contradictions: The GT context includes the original text, changed text, and contradicted text; b) Legal Contradictions: The context is augmented to include the original text, changed text, contradicted text, plus the specific contradicted law and a law explanation.
@@
For both location_alignment & law_match we have used GPT 4o and Gemini-2.5-Flash as judges.
@@
We have evaluated four language models: GPT-4o-mini, Gemini 2.0 Flash Team et al. (2025), Gemini 2.5 FlashComanici et al. (2025), and Llama 3.3 70b InstructGrattafiori et al. (2024). For consistency and reproducibility, all models were evaluated with standardized generation parameters. Hyperparameters mentioned in Appendix.
@@ C N C N C N C N
Table 7: Average Law_Match percentage averaged across categories where datasets (CUAD, NLI), levels (L1, L2), and judges (GPT4o vs Gemini 2.5).
@@ 6 Comparison with Related Works
Recent legal NLP advances have been propelled by datasets like CUAD Hendrycks et al. (2021) and LEDGAR Tuggener et al. (2020), which focus on clause labeling and categorization, as well as LexGLUE Chalkidis et al. (2022), ContractNLI Koreeda and Manning (2021), and CaseHOLD Zhong et al. (2020), which target tasks such as entailment and case retrieval. However, these resources rely on manual annotation and do not assess model reasoning under targeted perturbations. While recent work explores hallucination in legal text Shen et al. (2023) and LegalBench Zheng et al. (2023) broadens task diversity, they do not simulate document-level inconsistencies or robustness to semantic edits. Retrieval-augmented generation (RAG) has shown promise in open-domain QA Lin et al. (2022); Ji et al. (2023), but is rarely applied to legal contradiction detection.
CLAUSE is among the first fully automated, AI-driven legal benchmarks, using LLMs for both generation and validation. This scalable pipeline enables the creation of high-quality contradiction data, moving beyond static classification to simulate and verify realistic legal edits. Our approach uniquely bridges gaps in legal QA and hallucination detection, demonstrating the potential of automated dataset creation for complex legal reasoning tasks.
@@ 7 Conclusion
In this work, we present CLAUSE, a benchmark designed to evaluate legal reasoning in large language models via systematic contradiction detection. By combining persona-driven generation with RAG-based validation, we show that automated methods can produce high-quality, legally-grounded datasets. Our evaluations highlight significant gaps in current LLMs’ abilities to detect and explain nuanced contractual inconsistencies, revealing key challenges for legal AI. CLAUSE thus offers a foundation for advancing and validating legal AI systems, supporting the development of more reliable automated legal analysis tools. Through the four research questions explored in our experimental evaluation, covering performance, prompt sensitivity, generalization, and structured reliability, this study provides empirical evidence addressing the broader inquiry of whether current AI models are dependable enough for legal deployment. Our findings make this clear: despite impressive linguistic fluency, today’s LLMs lack the depth of legal reasoning and consistency required for high-stakes legal practice.
@@ A.2 Implementation Details and Package Documentation
A.3 Data Generation Prompts and Validation
@@ C.1 Evaluation 2: Miss and Extra Rates Across Categories
Category
GPT-4o mini Gemini 2.0 Gemini 2.5 Llama 3.3

### v2

@@ Abstract
Manan Roy Choudhury1thanks: The first three authors are recognized for their equal contribution to this work., Adithya Chandramouli [^1] footnotemark: 1, Mannan Anand [^1] footnotemark: 1, Vivek Gupta1 1Arizona State University {mroycho1, achand71, mannan.anand, vgupt140}@asu.edu
@@ 1 Introduction
The integration of Large Language Models (LLMs) into high-stakes legal work has outpaced the development of benchmarks that can audit their reliability against nuanced contractual flaws. This gap is critical, as a single undetected defect can have catastrophic consequences. In Perini Corp. v. Greate Bay Hotel & Casino, Inc., for instance, the simple omission of a consequential damages waiver resulted in a \$14.5 million liability, a sum over twenty times the contract’s fee McElroy, Deutsch, Mulvaney & Carpenter, LLP (2021). State-of-the-art models, optimized for fluency, are not trained to detect such legally significant risks, particularly those arising from the absence of text. Existing evaluation methods fail to probe these fragile reasoning capabilities Hu and others (2025); Mishra and others (2025), creating an urgent need for a new class of benchmark to stress-test LLMs against the adversarial and context-aware challenges of legal practice. This research directly confronts the ongoing debate about AI’s readiness for legal integration. By evaluating modern LLMs against our CLAUSE benchmark, we aim to provide empirical data to help answer the fundamental question: are these models truly reliable enough for such high-stakes work?
@@ 2 Our CLAUSE Dataset
Our pipeline applies a unified perturbation framework to the contracts. We extract full text and structural metadata (clause boundaries, section headers, cross-references) to preserve document integrity during modification. Contextually appropriate contradiction locations are identified using prompts across ten perturbation types. Three complementary prompting strategies ensure legal plausibility: persona-based roles (compliance officers, legal editors), few-shot examples, and chain-of-thought reasoning. Each perturbation produces a modified contract with structured JSON metadata containing original/altered text, perturbation type, document location, and model-generated explanations. Legal contradictions include law citations and statutory context; in-text contradictions include conflicting clause references and logical justifications. The unified process is applied to both datasets ensuring consistent output quality and metadata format, supporting downstream detection accuracy and legal reasoning analysis tasks.
@@
Table 1 presents comprehensive statistics for both components, showing the distribution of JSON files and perturbations across all ten categories. The dataset exhibits several notable characteristics. First, each JSON file contains multiple perturbations, with CUAD averaging 2.73 perturbations per file and the NLI split averaging 3.96 perturbations per file. Second, the distribution across categories is relatively balanced, though some variation exists due to the inherent complexity of generating specific perturbation types for different contract structures. Third, the legal categories generally contain fewer perturbations than their inText counterparts, reflecting the additional constraints imposed by statutory validation requirements. Extended data statistics for text length before and after perturbation is given in Table 10.
@@
To validate quality and legal accuracy, we first engaged a legal expert to confirm the real-world relevance of our perturbation categories. Subsequently, 3 NLP experts evaluated 25% of our dataset (13,275 perturbation instances) using structured binary classification to determine meaningful legal contradictions. Our human verification demonstrates exceptional quality with 98.58% contradiction validation rate and strong inter-rater reliability shown in Table 2. The evaluation examined semantic significance, contradiction strength, and contextual coherence, ensuring our benchmark contains high-quality examples for testing LLM legal reasoning capabilities. Legal expert verification statsin appendix Table 9.
@@ 5 Experimental Evaluation
explanation_match. The LLM, acting as a legal analyst, receives the combined explanations and all relevant GT contexts. It returns a structured JSON object with scores between 0-5 for: Accuracy (factual correctness), Completeness (coverage of key points), Clarity (ease of understanding), Legal Reasoning (soundness of legal logic), and a final binary Match flag that determines the alignment outcome. This process is adapted for two contradiction types. Ground-truth context are provided during the judging process for better context aware outputs: a) In-Text Contradictions: The GT context includes the original text, changed text, and contradicted text; b) Legal Contradictions: The context is augmented to include the original text, changed text, contradicted text, plus the specific contradicted law and a law explanation. Comparison between LLM-as-judge and human judge is in appendix Table 8.
@@
For both explanation_match & law_match we have used GPT 4o and Gemini-2.5-Flash as judges.
@@
We have evaluated four language models: GPT-4o-mini, Gemini 2.0 Flash Team et al. (2025), Gemini 2.5 FlashComanici et al. (2025), and LLaMa 3.3 70b InstructGrattafiori et al. (2024). For consistency and reproducibility, all models were evaluated with standardized parameters. Hyperparameters mentioned in Appendix Section A.
@@ C N C N C N C N
Table 7: Average Law_Match percentage averaged across categories where datasets (CUAD, NLI), levels (L1, L2), and judges (GPT4o vs Gemini 2.5 Flash).
@@ 6 Comparison with Related Works
Recent legal NLP advances have been propelled by datasets like CUAD Hendrycks et al. (2021) and LEDGAR Tuggener et al. (2020), which focus on clause labeling and categorization, as well as LexGLUE Chalkidis et al. (2022), ContractNLI Koreeda and Manning (2021), and CaseHOLD Zhong et al. (2020), which target tasks such as entailment and case retrieval. However, these resources rely on manual annotation and do not assess model reasoning under targeted perturbations. While recent work explores hallucination in legal text Shen et al. (2023) and LegalBench Zheng et al. (2023) broadens task diversity, they do not simulate document-level inconsistencies or robustness to semantic edits. Retrieval-augmented generation (RAG) has shown promise in open-domain QA Lin et al. (2022); Ji et al. (2023), but is rarely applied to legal contradiction detection. CLAUSE is among the first fully automated, AI-driven legal benchmarks, using LLMs for both generation and validation. This scalable pipeline enables the creation of high-quality contradiction data, moving beyond static classification to simulate and verify realistic legal edits. Our approach uniquely bridges gaps in legal QA and hallucination detection, demonstrating the potential of automated dataset creation for complex legal reasoning tasks.
@@ 7 Conclusion
In this work, we present CLAUSE, a benchmark designed to evaluate legal reasoning in large language models via systematic contradiction detection. By combining persona-driven generation with RAG-based validation, we show that automated methods can produce high-quality, legally-grounded datasets. Our evaluations highlight significant gaps in current LLMs’ abilities to detect and explain nuanced contractual inconsistencies, revealing key challenges for legal AI. CLAUSE thus offers a foundation for advancing and validating legal AI systems, supporting the development of more reliable automated legal analysis tools. Through the four research questions explored in our experimental evaluation, covering performance, prompt sensitivity, generalization, and structured reliability, this study provides empirical evidence addressing the broader inquiry of whether current AI models are dependable enough for legal deployment. Our findings make this clear: despite impressive linguistic fluency, today’s LLMs lack the depth of legal reasoning required for high-stakes legal practice.
Acknowledgements
We are grateful to Rui Heng Foo for his invaluable assistance with the data generation pipelines. We extend our sincere thanks to Arizona State University’s Lincoln Center for Applied Ethics, along with their law professors and partners, for lending their expertise to advise on our perturbation categories and to authenticate our perturbed contracts for real-world relevance. We also thank the ASU Law and Pre-Law students for volunteering their time and expertise. Finally, we acknowledge the CoRAL Lab at Arizona State University for providing the computational resources that made this research possible.
@@ A.2 Implementation Details and Package Documentation
A.3 Human Validation and Legal Expert Annotation
LLM-as-judge vs. Human Expert. To see whether the LLM-as-judge truly applies the defined rubrics correctly and how this aligns with human expert judgment, we have conducted a comparison study between LLM judges’ outputs against human-curated ground-truth labels using a parallel evaluation process on a sample (15% of the output sample) of the output generated. We used the human-validated ground-truth JSON file as the reference for scoring. Each file contains the corresponding ground-truth explanation, along with metadata linking them to specific discrepancy categories. Example snippet of such a json file is given at the end of the paper.
So, we use this human validated ground truth file explanation along with the other metadata and the predicted explanations and humans score the predicted explanations on the 4 parameters. The human experts assign the scores with all prior knowledge about the parameter definitions and the task. Here, legal experts are not required because we have to just match two different explanations and scoring them based on 4 different parameters which is an ideal task for NLP experts to take on.
Now below is the table 8 where we have reported the average difference of human and LLM as a judge score for the 4 parameters for a sample of the predicted results:
Metric LLM-as-Judge Expert Diff.
Accuracy 3.12 3.28 +0.16
Completeness 2.52 2.72 +0.20
Clarity 4.04 3.76 0.28-0.28
Legal Reasoning 3.16 2.92 0.24-0.24
Table 8: Comparison between LLM-as-Judge and Human Expert evaluations. All the reported values are mean values.
Legal Expert Validation. We engaged a licensed legal practitioner to audit approximately 10 perturbed documents from each of the 10 discrepancy categories ($\approx$100 total). Due to financial constraints, we were unable to extend the legal verification to a larger portion of the dataset. The expert reviewed these samples independently and confirmed that the perturbations were legally plausible, contextually coherent, and representative of real-world contractual risk patterns such as omissions, ambiguities, and misaligned terminology. The legal expert evaluated the perturbed documents based on the following four primary parameters:
Legal Plausibility (LP) – Whether the perturbation represented a legally coherent and realistic modification that could occur in actual contract drafting or review scenarios.
Contextual Coherence (CC) – Whether the changed clause remained semantically and contextually consistent within the overall document structure.
Contradiction Strength and Significance (CSS) – Whether the modification created a clear, unambiguous conflict, either internal (in-text) or external (statutory), that could plausibly affect compliance.
Representativeness of Legal Risk Type (RLR) – Whether the perturbation accurately reflected real-world contractual risk patterns such as omissions, ambiguities, misaligned terminology, inconsistencies, or structural flaws.
The expert rated each perturbation on a 0–5 scale across four above mentioned key parameters: 0 = Not acceptable, 5 = Fully satisfactory. The scores shown in Table 9 represent the averaged ratings across the reviewed samples.
Parameter Contradiction Type Expert Score
LP Outer-law Contradiction 4.83
LP In-text Contradiction 4.89
$C$ Outer-law Contradiction 4.56
$C$ In-text Contradiction 4.43
CSS Outer-law Contradiction 4.49
CSS In-text Contradiction 4.54
RLR Outer-law Contradiction 4.79
RLR In-text Contradiction 4.85
Table 9: Average expert evaluation scores (0–5) across contradiction types. Expert scores are averaged out. The abbreviations used are Legal Plausibility (LP), Contextual Coherence ($C$), Contradiction Strength and Significance (CSS), and Representativeness of Legal Risk Type (RLR).
A.4 Extended Data Statistics
In this sub-section we provide a detailed overview of the average text lenth before and after perturbation. We computed the character lengths of each original_text and changed_text entry across all ground-truth JSON files. The cumulative statistics for each dataset variant are summarized below in Table 10.
Dataset Cat Org. TL Changed TL Diff.
CUAD In-Text 546.2 496.7 49.5
CUAD Legal 510.7 503.8 6.90
NLI In-Text 470.5 414.0 56.5
NLI Legal 431.6 387.3 44.3
Table 10: Average text length statistics before and after perturbation. TL denotes text length.
A.5 Data Generation Prompts and Validation
@@ C.1 Evaluation 2: Miss and Extra Rates Across Categories
Category GPT-4o mini Gemini 2.0 Gemini 2.5 LLaMa 3.3